# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v2

@@ VI.4 Time operator
The results of this section demonstrate that the $(t,\phi)$ formalism is not a mere reparametrization of the periodic time $t^{\prime}$ from the $(t,t^{\prime})$ formalism, but that it is instead essential to correctly identify the thermodynamic role of the Floquet charge $F=Q_{\phi}$ and to compute the dynamics of the temporal FNG mode $z_{\phi}$. In particular, we can explicitly write the time evolution of its quantum amplitude,
@@
implies that, if $N$ is an eigenvalue of $\hat{N}$, then $N-N_{0}$ is also. This violates two fundamental properties of the spectrum of the particle number operator: its positive definiteness and its discreteness. Nevertheless, one can define phase fluctuations in situations involving large particle numbers, where one can neglect the discreteness of the spectrum and states with low occupation number, as in the case of a condensate. Analogously, by identifying energy with particle number and time with phase, one can define time fluctuations for large energies, well above the ground state. Actually, within the $(t,\phi)$ formalism, the time operator behaves as a sort of phase operator. This analogy suggests that spontaneous Floquet states should be highly excited states, shifted by a macroscopic energy from the ground state.
@@ VII Time supersolids: CES state
Due to the phase and time-translational invariance, if $\Psi_{0}(x,t)$ is a solution of Eq. (154), then $e^{-i\theta}\Psi_{0}(x,t+t_{0})$ is also a solution. In the $(t,\phi)$ formalism, where $\Psi_{0}(x,t_{0}+t)=\Psi_{0}(x,\phi_{0}+\omega t)$, the CES state is characterized by the vectors $\alpha=(\theta,t)$, $T_{\alpha}=(1,i\partial_{\phi})$, $Q_{\alpha}=(N,F)$, and $\lambda^{\alpha}=(\mu,\omega)$. The Floquet-Goldstone-Gibbs modes are obtained as
@@
The GS/CES phase diagram is an example of dynamical phase transition Moeckel and Kehrein (2008); Sciolla and Biroli (2010); Lang et al. (2018), where the GS is the symmetry-unbroken phase, with continuous time translation symmetry, while the CES state is the time-crystalline phase, with discrete time translation symmetry. The oscillation frequency $\omega$ exhibits a critical behavior close to the phase boundary, where the critical exponents for $v,Z$ (obtained from a fit in Fig. 1d) are both approximately $\alpha\simeq\beta\simeq 0.50$, in agreement with those for a square well de Nova and Sols (2022), potentially suggesting a possible analytical derivation.
@@
Figure 4: Normalized density-density correlation function $G(x,-x,t)$ for $x=10$ (left column), $x=20$ (center column), and $x=30$ (right column). Solid blue is the numerical result from a Truncated Wigner simulation and dashed red is the contribution from the temporal FNG mode, $G(x,x^{\prime},t)=r_{t}(x,t)r_{t}(x^{\prime},t)A(t)$, where $A(t)$ is numerically fitted as in Fig. 3 and $r_{t}(x,t)$ is computed from the time derivative of the actual CES wavefunction. Upper row: $Z=1$ and $v=0.8$. Lower row: $Z=1$ and $v=0.7$.
@@
$\delta\hat{n}(x,t)\simeq\Psi_{0}^{*}(x,t)\hat{\varphi}(x,t)+\Psi_{0}(x,t)\hat{\varphi}^{\dagger}(x,t)=iz^{\dagger}_{\theta}\sigma_{z}\hat{\Phi}.$ (168)
@@
$+$ $\sum_{\nu,\varepsilon}\hat{\gamma}_{\nu,\varepsilon}r_{\nu,\varepsilon}(x,t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\nu,\varepsilon}r^{*}_{\nu,\varepsilon}(x,t)e^{i\varepsilon t},$
$r_{n}(x,t)=iz^{\dagger}_{\theta}(x,t)\sigma_{z}z_{n}(x,t)$ being the density amplitude associated to the spinor $z_{n}$. As physically expected, the FNG mode associated to the phase does not couple to the density, $r_{\theta}=0$. Thus, only the temporal FNG mode $z_{t}(x,t)$ contributes to the correlation function $G(x,x^{\prime},t)$. Moreover, its amplitude $\hat{X}^{t}(t)=\hat{X}^{\phi}(t)/\omega$ grows linearly in time according to Eq. (149), so it eventually dominates the correlations,
@@
Due to the ballistic evolution of $\hat{X}^{t}(t)$, we expect a quadratic polynomial dependence for $A(t)$, $A(t)=at^{2}+bt+c$. This fits well the observed growth, black lines in upper row of Fig. 3. A more detailed comparison is presented in Fig. 4, where we directly compare the numerical results for $G(x,x^{\prime},t)$ from the Truncated Wigner method (solid blue) with the theoretical prediction of Eq. (170), combining the previous quadratic fit of $A(t)$ with a computation of $r_{t}(x,t)$ from the time derivative of the mean-field CES wavefunction itself (dashed red). A remarkable agreement is observed, confirming the ballistic motion of the quantum amplitude of the temporal FNG mode as the mechanism responsible for the quasi-periodic correlations.
@@
which requires the use of non-destructive imaging in order to measure the density at two consecutive times $t>t^{\prime}$, as in Ref. Smits et al. (2018). This caveat is not needed for the usual $U(1)$ symmetry breaking, where each realization of the experiment is insensitive to the global condensate phase because observables always commute with the total particle number, the generator of global phase transformations.
@@ VIII Discussion and perspectives
The Nambu-Goldstone modes associated to each spontaneously broken symmetry emerge as zero-energy modes of the Bogoliubov spectrum. Each NG mode is paired with a Gibbs mode, describing the fluctuations of the condensate wavefunction with respect to the associated conserved charge. These fluctuations arise because a symmetry-broken state cannot be an eigenstate of the charge operator that generates the corresponding broken-symmetry transformation. The orthogonalization of the Goldstone-Gibbs sector involves the Berry-Gibbs connection, a Berry connection whose variables are the conserved charges, which are not intrinsic parameters of the Hamiltonian. In contrast to the standard case, the Berry-Gibbs connection is not invariant under generalized gauge transformations, which expands the usual notion of phase transformations to include more general symmetry transformations. When extended to the complete manifold that includes the continuous parameters of the spontaneously broken symmetries, the Berry-Gibbs connection gives rise to a symplectic structure that allows to quantize the amplitudes of the Nambu-Goldstone/Gibbs modes as conjugate coordinate/momentum operators. This analogy is translated to the dynamics, where the momenta are time-independent as they represent the quantum fluctuations of the corresponding conserved charges, while the coordinates display a ballistic motion with velocities proportional to the momenta. The physical intuition behind this dynamics is very simple: since there is no restoring force along the orbits generated by the broken-symmetry transformations, fluctuations in those directions, spanned by the NG modes, lead to unbounded motion with constant velocity. Finally, we show that the Goldstone-Gibbs sector can be diagonalized, leading to a new set of independent NG modes. Remarkably, this diagonalization only involves thermodynamic considerations, and describes how the original NG modes hybridize within the different Bogoliubov branches in the low-frequency limit.
@@
The results of Secs. III, IV are thus of interest for the study of states which simultaneously break several symmetries, such as supersolids, quantum droplets, spinor and rotating condensates, or quantum solitons. They can also be easily translated to magnonic condensates Mäkinen et al. (2023), nonlinear optical fibers Drummond and Hillery (2014), and quantum fluids of light Carusotto and Ciuti (2013), as they obey nonlinear equations of motion similar to the Gross-Pitaevskii equation. Indeed, the generality of the formalism should allow for a straightforward adaptation to different systems, including fermions. For instance, within the $\nu=0$ quantum Hall phase diagram in graphene, states with two spontaneously broken symmetries were predicted within a Hartree-Fock description de Nova and Zapata (2017), of which some experimental evidence was recently reported Coissard et al. (2022). Indeed, Ref. de Nova and Zapata (2017) established the correspondence between the BdG and TDHFA equations, and translated Goldstone theorem into the specific context of the Hartree-Fock equations. Therefore, the BdG/TDHFA correspondence provides a direct link through which our results can be exported.
Another interesting perspective is provided by the geometry of the extended Berry-Gibbs connection. In this work, we have assumed that all the generators of the spontaneously broken symmetries commute between themselves. This leads, after a generalized gauge transformation, to a “flat” symplectic connection $F_{ab}=-\Omega_{ab}$. Actually, a generalized gauge transformation is nothing else than a change of coordinates in the extended Berry-Gibbs connection. However, if the generators of the spontaneously broken symmetries form a non-abelian Lie algebra, this no longer holds, opening the door for deeper geometrical implications, including nontrivial topological aspects. The exploration of this promising avenue is left for future work.
@@
In order to correctly describe the dynamics of the FNG modes, we develop the $(t,\phi)$ formalism, which goes beyond the usual $(t,t^{\prime})$ formalism. In combination with the generalized Gibbs ensemble, we provide a thermodynamical description of Floquet states analogous to that of stationary states, which we refer to as Floquet thermodynamics. This is possible because spontaneous Floquet states also have a well-defined energy, which enables the existence of a conserved Floquet charge. The Floquet charge is another hallmark of a spontaneous Floquet state, which adds to the presence of a temporal Floquet-Nambu-Goldstone mode. In fact, both features are the sides of the same coin, since they arise, via Goldstone and Noether theorems, from the continuous time-translation symmetry of the underlying time-independent Hamiltonian. Conventional Floquet states can also be described within the framework of Floquet thermodynamics by means of the Floquet enthalpy, i.e., the Legendre transform of the energy with respect to the Floquet charge, whose conjugated variable is in turn the frequency, fixed here by the external driving. By drawing an analogy with standard Thermodynamics, spontaneous Floquet states are “isofloquetic”, and driven Floquet states are “isoperiodic”.
@@
Finally, we apply our formalism to a specific realization of spontaneous Floquet state, the CES state, which spontaneously breaks $U(1)$ and continuous time-translation symmetry, behaving as a time supersolid. We propose an experimental scheme, based on standard analogue gravity techniques Steinhauer (2016); de Nova et al. (2019); Kolobov et al. (2021), to observe the temporal FNG mode, which does couple to the density in contrast to that of the phase. In particular, we prove that the density-density correlations are dominated at long times by the ballistic growth of the temporal FNG mode, from where it can be extracted. Numerical results from a Truncated Wigner simulation show a remarkable agreement with our theory.
@@
The results of Secs. V-VI can be used to describe simultaneous symmetry breaking in both spontaneous and conventional Floquet states. In particular, they establish a fundamental distinction between both states at the thermodynamical level, leading to an important conclusion: spontaneous Floquet states do not experience heating as they conserve energy, in contrast to conventional Floquet systems D’Alessio and Rigol (2014); Lazarides et al. (2014). In the specific case of the CES state, this survives for thermodynamically long times which scale linearly with the system size de Nova and Sols (2022). This can be compared with the prethermal regime of a driven Floquet system, which scales exponentially with the applied frequency Pizzi et al. (2021); Ye et al. (2021); Kyprianidis et al. (2021).
@@
The quest for a time operator is a fundamental question in Quantum Mechanics Aharonov and Bohm (1961); Susskind and Glogower (1964); Unruh and Wald (1989), which is still open Höhn et al. (2021). In our case, the time operator emerges from the quantum fluctuations of the global time origin, in close analogy with the fluctuations of the global phase of a condensate. Indeed, in the $(t,\phi)$ formalism, the time operator emerges from the fluctuations of the phase shift $\phi_{0}$ in the periodic motion of the spontaneous Floquet state. The connection between time and phase has been pointed out since early approaches to the problem Susskind and Glogower (1964) because any realization of a time operator is expected to be based on the oscillation of a clock. The time-phase analogy suggests that spontaneous Floquet states should be then macroscopically excited states, with an energy well above the ground state. As a result, the CES state provides a condensed-matter setup where a tangible realization of a time operator can be studied.
@@
Our work also finds a major application in the field of time crystals, since a spontaneous Floquet state represents a specific realization of a continuous time crystal de Nova and Sols (2022). Indeed, some results of our work are connected with the time crystal literature. For instance, the role of the generalized Gibbs ensemble in dissipative time crystals was thoroughly discussed in Ref. Booker et al. (2020). The emergence of continuous time crystals from excited eigenstates, in the same fashion as spontaneous Floquet states, was predicted in Ref. Syrwid et al. (2017). Time supersolids, like the CES state, were originally observed in Ref. Autti et al. (2018). This implies that the formalism of this work is directly applicable to time supersolids. On the other hand, Floquet thermodynamics can provide a novel tool for the characterization of both continuous and discrete time crystals. Moreover, the analysis presented in Sec. IV suggests that certain periodic systems identified as time crystals may not genuinely break continuous time-translation symmetry since they can arise from constant motion along the closed orbit generated by a broken-symmetry transformation. Thus, the presence of a genuine temporal FNG mode should be regarded as the characteristic signature of any bona fide continuous time crystal.
@@
Although we have focused on atomic condensates, our Floquet formalism can be easily extrapolated to magnonic condensates, nonlinear optical fibers or quantum fluids of light. In general, spontaneous Floquet states extend the field of nonlinear Floquet waves Kreil et al. (2019); Träger et al. (2021) to scenarios without external driving. Spontaneous Floquet states in fermionic systems described by HF approaches, such as quantum Hall states, are also plausible de Nova and Sols (2022). In addition, self-consistent BCS-type theories are known to display periodic oscillations in momentum space Barankov et al. (2004); Foster et al. (2014); Perfetto and Stefanucci (2020).

### v3

@@ VI.4 Time operator
The previous results demonstrate that the $(t,\phi)$ formalism is not a mere reparametrization of the periodic time $t^{\prime}$ from the $(t,t^{\prime})$ formalism, but that it is instead essential to correctly identify the thermodynamical role of the Floquet charge $F=Q_{\phi}$ and to compute the dynamics of the temporal FNG mode $z_{\phi}$. In particular, we can explicitly write the time evolution of its quantum amplitude,
@@
implies that, if $N$ is an eigenvalue of $\hat{N}$, then $N-N_{0}$ is also. This violates two fundamental properties of the spectrum of the particle number operator: its positive definiteness and its discreteness. Nevertheless, one can define phase fluctuations in situations involving large particle numbers, where one can neglect the discreteness of the spectrum and states with low occupation number, as in the case of a condensate. Analogously, by identifying energy with particle number and time with phase, one can define time fluctuations for large energies, well above the ground state. Actually, within the $(t,\phi)$ formalism, the time operator is a sort of phase operator. This analogy suggests that spontaneous Floquet states should be highly excited states, shifted by a macroscopic energy from the ground state.
@@ VII Time supersolids: CES state
Due to the phase and time-translational invariance, if $\Psi_{0}(x,t)$ is a solution of Eq. (156), then $e^{-i\theta}\Psi_{0}(x,t+t_{0})$ is also a solution. In the $(t,\phi)$ formalism, where $\Psi_{0}(x,t+t_{0})=\Psi_{0}(x,\omega t+\phi_{0})$, the CES state is characterized by the vectors $\alpha=(\theta,\phi)$, $T_{\alpha}=(1,i\partial_{\phi})$, $Q_{\alpha}=(N,F)$, and $\lambda^{\alpha}=(\mu,\omega)$. The Floquet-Goldstone-Gibbs modes are then
@@
The GS/CES phase diagram is an example of dynamical phase transition Moeckel and Kehrein (2008); Sciolla and Biroli (2010); Lang et al. (2018), where the GS is the symmetry-unbroken phase, with continuous time translation symmetry, while the CES state is the time-crystalline phase, with discrete time translation symmetry. The oscillation frequency $\omega$ exhibits a critical behavior close to the phase boundary, where the critical exponents for $v,Z$ (obtained from a fit in Fig. 2d) are both approximately $\gamma\simeq\delta\simeq 0.50$, in agreement with those for a square well de Nova and Sols (2022), potentially suggesting a possible analytical derivation.
@@
Figure 5: Normalized density-density correlation function $G(x,-x,t)$ for $x=10$ (left column), $x=20$ (center column), and $x=30$ (right column). Solid blue is the numerical result from a Truncated Wigner simulation and dashed red is the contribution from the temporal FNG mode, $G(x,x^{\prime},t)=r_{t}(x,t)r_{t}(x^{\prime},t)A(t)$, where $A(t)$ is numerically fitted as in Fig. 4 and $r_{t}(x,t)=\partial_{t}n(x,t)$ is computed from the time derivative of the mean-field density. Upper row: $Z=1$ and $v=0.8$. Lower row: $Z=1$ and $v=0.7$.
@@
$\delta\hat{n}(x,t)\simeq\Psi_{0}^{*}(x,t)\hat{\varphi}(x,t)+\Psi_{0}(x,t)\hat{\varphi}^{\dagger}(x,t)=-iz^{\dagger}_{\theta}\sigma_{z}\hat{\Phi}.$ (170)
@@
$+$ $\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}r_{\varepsilon,\nu}(x,t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}r^{*}_{\varepsilon,\nu}(x,t)e^{i\varepsilon t},$
$r_{n}(x,t)=-iz^{\dagger}_{\theta}(x,t)\sigma_{z}z_{n}(x,t)$ being the density amplitude associated to the spinor $z_{n}$. In the case of the Goldstone-Gibbs modes, the density amplitude is just
$r_{a}(x,t)=\partial_{a}|\Psi_{0}(x,t)|^{2}=\partial_{a}n(x,t),$ (172)
where $n(x,t)$ is the mean-field density, which is approximately equal to the ensemble-averaged density after neglecting higher-order corrections in the field fluctuations, $\braket{\hat{n}(x,t)}\simeq n(x,t)$.
As physically expected, the FNG mode associated to the phase does not couple to the density, $r_{\theta}=0$. Thus, only the temporal FNG mode $z_{t}(x,t)$ contributes to the correlation function $G(x,x^{\prime},t)$. Moreover, its amplitude $\hat{X}^{t}(t)=\hat{X}^{\phi}(t)/\omega$ grows linearly in time according to Eq. (151), so it eventually dominates the correlations,
@@
Due to the ballistic evolution of $\hat{X}^{t}(t)$, we expect a quadratic polynomial dependence for $A(t)$, $A(t)=at^{2}+bt+c$. This fits well the observed growth, black lines in upper row of Fig. 4. A more detailed comparison is presented in Fig. 5, where we depict together the numerical results for $G(x,x^{\prime},t)$ from the Truncated Wigner method (solid blue), and the theoretical prediction of Eq. (173), combining the previous quadratic fit of $A(t)$ with a computation of $r_{t}(x,t)=\partial_{t}n(x,t)$ from the time derivative of the mean-field density (dashed red). A remarkable agreement is observed, confirming the ballistic motion of the quantum amplitude of the temporal FNG mode as the mechanism responsible for the quasi-periodic correlations.
@@
which requires the use of non-destructive imaging in order to measure the density at two consecutive times $t>t^{\prime}$, as in Ref. Smits et al. (2018). This caveat is not needed for the measurement of correlations under the usual $U(1)$ symmetry breaking, where each realization of the experiment is insensitive to the global condensate phase because observables always commute with the total particle number, the generator of global phase transformations.
@@ VIII Discussion and perspectives
The Nambu-Goldstone modes associated to each spontaneously broken symmetry emerge as zero-energy modes of the excitation spectrum. Each NG mode is paired with a Gibbs mode, describing the fluctuations of the quantum state with respect to the associated conserved charge. These fluctuations arise because a symmetry-broken state cannot be an eigenstate of the charge operator that generates the corresponding broken-symmetry transformation. The orthogonalization of the Goldstone-Gibbs sector involves the Berry-Gibbs connection, a Berry connection whose variables are the conserved charges, which are not intrinsic parameters of the Hamiltonian. In contrast to the standard case, the resulting Berry-Gibbs curvature is not invariant under generalized gauge transformations, which expands the usual notion of phase transformations to include more general symmetry transformations. When extended to the complete manifold that includes the continuous parameters of the spontaneously broken symmetries, the Berry-Gibbs curvature gives rise to a symplectic structure that allows to quantize the amplitudes of the Nambu-Goldstone/Gibbs modes as conjugate coordinate/momentum operators. This analogy is translated to the dynamics, where the momenta are time-independent as they represent the quantum fluctuations of the corresponding conserved charges, while the coordinates display a ballistic motion with velocities proportional to the momenta. The physical intuition behind this dynamics is very simple: since there is no restoring force along the orbits generated by the broken-symmetry transformations, fluctuations in those directions, spanned by the NG modes, lead to unbounded motion with constant velocity. Finally, we show that the Goldstone-Gibbs sector can be diagonalized, leading to a new set of independent NG modes. Remarkably, this diagonalization only involves thermodynamic considerations, and describes how the original NG modes hybridize within the different branches of the spectrum in the low-frequency limit.
@@
The results of Secs. III, IV are thus of interest for the study of states which simultaneously break several symmetries, such as supersolids, quantum droplets, spinor and rotating condensates, or quantum solitons. They can also be easily translated to magnonic condensates Mäkinen et al. (2023), nonlinear optical fibers Drummond and Hillery (2014), and quantum fluids of light Carusotto and Ciuti (2013), as they obey nonlinear equations of motion similar to the Gross-Pitaevskii equation. Indeed, the generality of the formalism, based on a Lagrangian description, allows for a straightforward adaptation to arbitrary many-body systems, including fermions. For instance, within the $\nu=0$ quantum Hall phase diagram in graphene, states with two spontaneously broken symmetries were predicted within a Hartree-Fock description de Nova and Zapata (2017), of which some experimental evidence was recently reported Coissard et al. (2022). Indeed, Ref. de Nova and Zapata (2017) established the correspondence between the BdG and TDHFA equations, and translated Goldstone theorem into the specific context of the Hartree-Fock equations. Therefore, the BdG/TDHFA correspondence provides a direct link through which our results can be exported.
Another interesting perspective is provided by the geometry of the extended Berry-Gibbs connection. In this work, we have assumed that all the generators of the spontaneously broken symmetries commute between themselves. This leads, after a generalized gauge transformation, to a “flat” symplectic curvature $F_{ab}=-\Omega_{ab}$. Actually, a generalized gauge transformation is nothing else than a change of coordinates in the extended Berry-Gibbs manifold. However, if the generators of the spontaneously broken symmetries form a non-abelian Lie algebra, this no longer holds, opening the door for deeper geometrical implications, including nontrivial topological aspects. The exploration of this promising avenue is left for future work.
@@
In order to correctly describe the dynamics of the FNG modes, we develop the $(t,\phi)$ formalism, which goes beyond the usual $(t,t^{\prime})$ formalism. In combination with the generalized Gibbs ensemble, we provide a thermodynamical description of Floquet states analogous to that of stationary states, which we refer to as Floquet thermodynamics. This is possible because spontaneous Floquet states also have a well-defined energy, which enables the existence of a conserved Floquet charge. The Floquet charge is another hallmark of a spontaneous Floquet state, which adds to the presence of a temporal Floquet-Nambu-Goldstone mode. In fact, both features are the sides of the same coin, since they arise, via Goldstone and Noether theorems, from the continuous time-translation symmetry of the underlying time-independent Hamiltonian.
Conventional Floquet states can also be described within the framework of Floquet thermodynamics by means of the Floquet enthalpy, i.e., the Legendre transform of the energy with respect to the Floquet charge, whose conjugated variable is in turn the frequency, fixed here by the external driving. By drawing an analogy with standard Thermodynamics, spontaneous Floquet states are “isofloquetic”, and driven Floquet states are “isoperiodic”.
@@
Finally, we apply our formalism to a specific realization of spontaneous Floquet state, the CES state, which spontaneously breaks $U(1)$ and continuous time-translation symmetry, representing a time supersolid. We propose an experimental scheme, based on standard analogue gravity techniques Steinhauer (2016); de Nova et al. (2019); Kolobov et al. (2021), to observe the temporal FNG mode, which does couple to the density in contrast to the FNG mode of the phase. In particular, we prove that the density-density correlations are dominated at long times by the ballistic growth of the temporal FNG mode, from where it can be extracted. Numerical results from a Truncated Wigner simulation show a remarkable agreement with our theory.
@@
The results of Secs. V-VI can be used to describe simultaneous symmetry breaking in both spontaneous and conventional Floquet states. Furthermore, the $(t,\phi)$ formalism and the Floquet thermodynamics can be applied to an arbitrary many-body system, where they establish a fundamental distinction between both states at the thermodynamical level, leading to an important conclusion: spontaneous Floquet states do not experience heating as they conserve energy, in contrast to conventional Floquet systems D’Alessio and Rigol (2014); Lazarides et al. (2014). In the specific case of the CES state, this survives for thermodynamically long times which scale linearly with the system size de Nova and Sols (2022). This can be compared with the prethermal regime of a driven Floquet system, whose lifetime scales exponentially with the applied frequency Pizzi et al. (2021); Ye et al. (2021); Kyprianidis et al. (2021).
@@
The quest for a time operator is a fundamental question in Quantum Mechanics Aharonov and Bohm (1961); Susskind and Glogower (1964); Unruh and Wald (1989), which is still open Höhn et al. (2021). In our case, the time operator emerges from the quantum fluctuations of the global time origin, in close analogy with the fluctuations of the global phase of a condensate. Indeed, in the $(t,\phi)$ formalism, the time operator emerges from the fluctuations of the global phase shift $\phi_{0}$ in the periodic motion of the spontaneous Floquet state. The connection between time and phase has been pointed out since early approaches to the problem Susskind and Glogower (1964) because any realization of a time operator is expected to be based on the oscillation of a clock. The time-phase analogy suggests that spontaneous Floquet states should be then macroscopically excited states, with an energy well above the ground state. As a result, the CES state provides a condensed-matter setup where a tangible realization of a time operator can be studied.
@@
Our work also finds a major application in the field of time crystals, since a spontaneous Floquet state represents a specific realization of a continuous time crystal de Nova and Sols (2022). Indeed, some results of our work are connected with the time-crystal literature. For instance, the role of the generalized Gibbs ensemble in dissipative time crystals was thoroughly discussed in Ref. Booker et al. (2020). The emergence of continuous time crystals from excited eigenstates, in the same fashion as spontaneous Floquet states, was predicted in Ref. Syrwid et al. (2017). Time supersolids, like the CES state, were originally observed in Ref. Autti et al. (2018), to which the formalism of this work is directly applicable. On the other hand, Floquet thermodynamics can provide a novel tool for the characterization of both continuous and discrete time crystals. Moreover, the analysis presented in Sec. IV suggests that certain periodic systems identified as time crystals may not genuinely break continuous time-translation symmetry since they can arise from constant motion along a closed orbit generated by the broken-symmetry transformations. Thus, the presence of a genuine temporal FNG mode should be regarded as the characteristic signature of any bona fide continuous time crystal.
@@
Although we have used atomic condensates to illustrate our Floquet formalism, it can be straightforwardly extrapolated to magnonic condensates, nonlinear optical fibers or quantum fluids of light. In general, spontaneous Floquet states extend the field of nonlinear Floquet waves Kreil et al. (2019); Träger et al. (2021) to scenarios without external driving. Spontaneous Floquet states in fermionic systems described by HF approaches, such as quantum Hall states, are also plausible de Nova and Sols (2022). In addition, self-consistent BCS-type theories are known to display periodic oscillations in momentum space Barankov et al. (2004); Foster et al. (2014); Perfetto and Stefanucci (2020).